OpenStudy (anonymous):

What is the 1/5x+1/3y=4/5 and 1/5x+2y=54/5 if using the method of substitution?

ganeshie8 (ganeshie8):

isolate 1/5x value from the first equation, substitute it in the second equation

OpenStudy (anonymous):

How do I do that?

ganeshie8 (ganeshie8):

Your first equation : 1/5x+1/3y=4/5 subtract 1/3y both sides, what do you get ?

OpenStudy (anonymous):

1/5x = 1/2?

ganeshie8 (ganeshie8):

careful, you cannot subtract apples and coconuts

ganeshie8 (ganeshie8):

substitute this value in second equation and solve y

OpenStudy (anonymous):

1/5x + 2(5-1/3y)=54/5

ganeshie8 (ganeshie8):

noo, you need to substitute it for 1/5x

OpenStudy (anonymous):

combine the y's?

ganeshie8 (ganeshie8):

you need to get the common denominator to combine fractions

OpenStudy (anonymous):

this is not looking right: 3/3y = 54

ganeshie8 (ganeshie8):

yeah we made a mistake, lets work it again

ganeshie8 (ganeshie8):

1/5x+1/3y=4/5 subtract 1/3y both sides : 1/5x = 4/5 - 1/3y substitute this value in second equation : 4/5 - 1/3y +2y=54/5 multiply through by 15 : 12 - 5y + 30y = 162 25y = 150 y = 6

ganeshie8 (ganeshie8):

substitute the y value in any one of the given equations and solve x

OpenStudy (anonymous):

x=-.59

ganeshie8 (ganeshie8):

that doesn't look correct... check your work once...

ganeshie8 (ganeshie8):

1/5x+1/3y=4/5 plugin y = 6 1/5x + 1/3(6) = 4/5 1/5x + 2 = 4/5 multiply through by 5 : x + 10 = 4 x = ?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

(-6,6)

ganeshie8 (ganeshie8):

Looks good !

OpenStudy (anonymous):

thank you!!!